A force of 9n acts along the positive x axis and a force of 4n acts in direction making an angle of 60⁰ with the positive direction of the x axis. What is the sum of the resolved forces along the x axis ?

1
Expert's answer
2022-02-18T08:22:30-0500

The sum of the resolved forces along the x axis

"F_x=F_{1x}+F_{2x}=F_1+F_2\cos\theta"

"F_x=\rm 9\: N+4\: N\cos 60^\circ=11\: N"